OpenStudy (anonymous):

Combine like terms to solve for the variable. check your answer 9h+3-5h-8+3h-5=3h+2-6h+8-5h-5 If the perimeter of a triangle is 94 and the sides are (3d+2) in., (2d+1) in. and (5d-9) in., find the following a. what is the value of d? b.what is the length of each side? Please help i got stuck on this two last question please

OpenStudy (anonymous):

9h+3-5h-8+3h-5=3h+2-6h+8-5h-5 9h-5h+3h-3h+6h+ 5h = -3+8+5+2+ 8-5 15h = 15 h = 1 checking LHs = 9+3-5-8+3-5 = -3 RHS = 3+2-6+8-5-5 = -3 OK

OpenStudy (anonymous):

first add um the "H" numbers and add up the normal numbers(no H)

OpenStudy (anonymous):

3d+2+2d+1+5d-9 = 94 10d = 94 - 2 -1 + 9 = 100 d = 10 length of sides = 32, 21 and 41